OpenStudy (anonymous):

Physics: Solve v=d/t for t when d=50, v=30

OpenStudy (anonymous):

First rearrange the equation for t in terms of v and d. then just plug in the given values...

OpenStudy (anonymous):

Just subs

OpenStudy (anonymous):

30 = 50/t 30t = 50 t = 50/30 t=5/3 = 1.66

OpenStudy (ash2326):

@Manita did you understand?

OpenStudy (anonymous):

yes thank you :)